# Generalizable Spectral Embedding with an Application to UMAP

### v1

@@ 4.1 Motivation
It is well known that the matrix $V\in\mathbb{R}^{n\times k}$, containing the first $k$ eigenvectors of $L$ (i.e., those corresponding to the $k$ smallest eigenvalues) as its columns, minimizes $R_{L}(A)$ under orthogonality constraint (i.e. $A^{T}A=I$) (Li, 2015).
@@
An immediate result of Lemma 1 is that SpectralNet’s method, using a deep neural network for RQ minimization (while enforcing orthogonality), does not lead to the SE. However, it only leads to the space spanned by the constant vector and the leading $k-1$ eigenvectors of $L$, with different rotations and reflections for each run. Therefore, each time the RQ is minimized, it results in a different linear combination of the smallest eigenvectors. Although this is sufficient for clustering purposes, as we search for reproducibility, consistency, and separation of the eigenvectors, the RQ cannot solely provide the SE, necessitating the development of new techniques in GrEASE.
4.2 GrEASE
@@
Here we present the two key components of GrEASE, a scalable and generalizable SE method. We consider the following setup: Given a training set $\mathcal{X}\subseteq\mathbb{R}^{d}$ and a target dimension $k$, we construct an affinity matrix $W$, and compute an approximation of the leading eigenvectors of its corresponding Laplacian. In practice, we first utilize SpectralNet (Shaham et al., 2018) to approximate the space spanned by the first $k+1$ eigenfunctions of the corresponding Laplace-Beltrami operator, and then find each of the $k$ leading eigenfunctions within this space (i.e. the SE). Namely, GrEASE computes a map $F_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{k}$, which approximates the map $\bar{f}=(f_{2},\dots,f_{k+1})$, where $f_{i}$ is the $i$th eigenfunction of the Laplace-Beltrami operator on the underlying manifold $\mathcal{D}$.
@@
As empirically showed in (Shaham et al., 2018), and motivated from Lemma 1, SpectralNet loss is minimized when $F_{\theta}=T\circ(f_{1},\dots,f_{k+1})$, where $T:\mathbb{R}^{k+1}\to\mathbb{R}^{k+1}$ is an arbitrary isometry. That is, $F_{\theta}$ approximates the space spanned by the first $k+1$ eigenfunctions. However, the SE (i.e. each of the leading eigenfunctions) is poorly approximated. Each time the RQ is minimized, the eigenfunctions are approximated up to a different isometry $T$. Fig. 2(a) demonstrates this phenomenon on the toy moon dataset - a noisy half circle linearly embedded into 10-dimension input space (see Sec. 5.1). Employing SpectralNet approach indeed enables us to consistently achieve a perfect approximation of the space (i.e., the errors at the left histograms are accumulated around 0). However, when comparing vector to vector, it becomes apparent that the SE was seldom attained. That is, the distances are spread across the entire range from 0 to 1.
@@
To get the SE consistently (i.e., to separate the eigenvectors), we suggest a simple use of Lemma 1. Notice that based on Lemma 1 we can compute a rotated version of the diagonal eigenvalues matrix. Namely,
$(VQ)^{T}L(VQ)=Q^{T}V^{T}LVQ=Q^{T}\Lambda Q=:\tilde{\Lambda}$
Where $\Lambda$ is the diagonal eigenvalues matrix. Due to the uniqueness of eigendecomposition, the eigenvectors and eigenvalues of the small matrix $\tilde{\Lambda}\in\mathbb{R}^{k+1\times k+1}$ are $Q^{T}$ and $diag(\Lambda)$, respectively. Hence, by diagonalizing $\tilde{\Lambda}$ we get the eigenvalues and are also able to separate the eigenvectors (i.e., approximate the SE).
@@
The effect of this intentional rotation is represented in the Fig. 2(a). GrEASE was not only able to consistently approximate the space, but also approximate each eigenvector. While SpectralNet errors are distributed over a large range of values, GrEASE errors are small, capturing only the smallest error bin in the figure.
Figure 2: (a) GrEASE separates the eigenvectors. Approximation of the 2-dimensional SE of the moon dataset using SpectralNet (in blue) and GrEASE (in green) over 100 runs, on train (top row) and test (bottom row). Left column: distribution of the Grassmann distance between the output and true subspace. Second to Fourth columns: distribution of the $sin^{2}$ distance between each output and true eigenvector separately. Evidently, GrEASE is able to separate the eigenvectors. (b) GrEASE is scalable. Running times of SE using GrEASE vs. other methods on the Moon dataset (a 2D moon linearly embedded into 10D input space), relative to the number of samples, and with standard deviation confidence intervals. Evidently, GrEASE is the fastest asymptotically.
@@
Our end-to-end training approach is summarized in Algorithms 1 and 2 in Appendix B. We run them consecutively: First, we train $F_{\theta}$ to approximate the first eigenfunctions up to isometry (Algorithm 1) (Shaham et al., 2018). Second, we find the matrices $Q^{T}$ and $\Lambda$ to separate the eigenvectors and retrieve the SE and its corresponding eigenvalues (Algorithm 2). App. C details additional considerations about the implementation.
@@
As the network iterates over small batches, and the post-processing operation is much cheaper, GrEASE’s time complexity is approximately linear in the number of samples. This is also demonstrated in Fig. 2(b), where the continuous red line, representing linear regression, aligns with our empirical results. App. C provides a discussion about the complexity of GrEASE. Note also that GrEASE is much more memory-efficient than existing methods, as it does not require storing the full graph, or any large matrix, in the memory, but rather one small graph or matrix (of a minibatch) at a time.
@@ 4.3 NUMAP
We focus on GrEASE application to UMAP, one of many methods which can benefit from a generalizable SE. As discussed in Sec. 3.3, the SE initialization is crucial for the global preservation abilities of UMAP. Therefore, we seek a method to incorporate SE into a generalizable version of UMAP. It is important to note that a naive approach would be to fine-tune GrEASE using UMAP loss. However, during this implementation, we encountered the phenomenon of catastrophic forgetting (see App. F).
The core of our idea is illustrated in Fig. 3. Initially, we use GrEASE to learn a parametric representation of the $k$-dimensional SE of the input data. Subsequently, we train an NN to map from the SE to the UMAP embedding space, utilizing UMAP contrastive loss. The objective of the second NN is to identify representations that best capture the local structure of the input data graph. SE transforms complex non-linear structures into simpler linear structures, allowing the second NN to preserve both local and global structures effectively. To enhance this capability, we incorporate residual connections from the first to the last layer of the second NN. Specifically, the objective is to minimize the residual between the $\ell$-dimensional UMAP embedding and the $\ell$-dimensional SE. Note that this could not have been made possible without GrEASE’s ability to separate the eigenvectors (and would not be practical without its inherent generalizability and scalability). Fig. 1a demonstrates this capability with several simple structures.
Figure 3: Incorporating GrEASE for generalizable UMAP. UMAP vs. NUMAP vs. P. UMAP overview. A green arrow represents a non-parametric step. NUMAP integrates SE, as in UMAP, while enabling generalization.
@@ 4.4 Additional Applications
In this section we seek to highlight GrEASE’s potential impact on important tasks and applications (besides UMAP), as it integrates generalizability, scalability and eigenvectors separation. As discussed in Sec. 1, SE is applied across various domains, many of which can benefit generalizability capabilities by simply replacing the current SE implementation with GrEASE. We therefore elaborate herein the significance of SE in selected applications, and discuss how GrEASE, as a generalizable approximation of it, can enhance their effectiveness and applicability.
@@
A special case of SE is the Fiedler vector and value (Fiedler, 1973, 1975). The Fiedler value, also known as algebraic connectivity, refers to the second eigenvalue of the Laplacian matrix, while the Fiedler vector refers to the associated eigenvector. This value quantifies the connectivity of a graph, increasing as the graph becomes more connected. Specifically, if a graph is not connected, its Fiedler value is 0. The Fiedler vector and value are a main topic of many works (Andersen et al., 2006; Barnard et al., 1993; Kundu et al., 2004; Shepherd et al., 2007; Cai et al., 2018; Zhu and Schlick, 2021; Tam and Dunson, 2020).
As GrEASE is able to distinguish between the eigenvectors and approximate the eigenvalues, it has the capability to approximate both the Fiedler vector and value, while also generalizing the vector to unseen samples (see Sec. 5.1).
@@
A popular method which incorporates SE, alongside the eigenvalues of the Laplacian matrix, is Diffusion Maps (Coifman and Lafon, 2006a). Diffusion Maps embeds a graph (or a manifold) into a space where the pairwise Euclidean distances are equivalent to the pairwise Diffusion distances on the graph.
@@
$X\to\begin{pmatrix}(1-\lambda_{2})^{t}v_{2}&\cdots&(1-\lambda_{k+1})^{t}v_{k+1}\end{pmatrix}=Y$
Where $X\in\mathbb{R}^{n\times d}$ is a matrix containing each input point as a row, and $Y\in\mathbb{R}^{n\times k}$ is a matrix containing each of the representations as a row. As GrEASE is able to approximate both the eigenvectors and eigenvalues of the Laplacian matrix, it is able to make Diffusion Maps generalizable and efficient (Sec. 5.1).
@@ 4.5 Evaluating UMAP embedding - Grassmann Score
To address this gap, we propose a new evaluation method, specifically appropriate for assessing global structure preservation in graph-based dimensionality reduction methods (e.g., UMAP, t-SNE). The leading eigenvectors of the Laplacian matrix are known to encode crucial global information about the graph (Belkin and Niyogi, 2003). Thus, we measure the distance between the global structures of the original and embedding manifolds using the Grassmann distance between the first eigenvectors of their respective Laplacian matrices. We refer to this method as the Grassmann Score (GS).
It is important to note that GS includes a hyper-parameter - the number of eigenvectors considered. Increasing the number of eigenvectors incorporates more local structure into the evaluation. A natural choice for this hyperparameter is 2, which corresponds to comparing the Fiedler vectors (i.e., the second eigenvectors of the Laplacian). The Fiedler vector is well known for encapsulating the global information of a graph (Fiedler, 1973, 1975). Unless stated otherwise, we use two eigenvectors for computing the GS. Fig. 4 demonstrates GS (alongside Silhouette and kNN scores for comparison) on a few embeddings of two tangent spheres, independently to the embedding methods. Notably, the embedding on the right appears to best preserve the global structure, as indicated by the smallest GS value. In contrast, the kNN scores are comparable across all embeddings (e.g., kNN ignores separation of an existing class), and the Silhouette score even favors other embeddings. In App. D we mathematically formalize GS and provide additional examples of embeddings and their corresponding GS. These examples further support the intuition that GS effectively captures global structure preservation better than previous measures.
Figure 4: Grassmann Score (GS) captures global structure preservation. A demonstration of the alignment between the intuitive expectation and the GS results on a toy dataset of two 3-dimensional tangent spheres. Four possible 2-dimensional embeddings of this dataset are provided, along with their corresponding GS, kNN accuracy and Silhouette score. Unlike kNN and Silhouette, GS effectively captures the preservation of global structure.
@@ 5.1 Eigenvectors Separation - Generalizable SE
In this section, we demonstrate GrEASE’s ability to approximate and generalize the SE using four real-world datasets: CIFAR10 (via their CLIP embedding); Appliances Energy Prediction dataset (Candanedo, 2017); Kuzushiji-MNIST (KMNIST) dataset (Clanuwat et al., 2018); Parkinsons Telemonitoring dataset (Tsanas and Little, 2009). Particularly, we compare our results with SpectralNet, which has been empirically shown to approximate the SE space. However, as our results demonstrate, SpectralNet is insufficient for accurately approximating SE. For additional technical details regarding the datasets, architectures and training procedures, we refer the reader to Appendix G.
@@
To assess the approximation of each eigenvector (i.e., the SE), we compute the $sin^{2}$ of the angle between each predicted and ground truth vector. This can be viewed as the 1-dimensional case of the Grassmann distance, a well-known metric for comparing equidimensional linear subspaces (see formalization in App. D). Concerning the eigenvalues approximation evaluation, we measure the Pearson Correlation between the predicted and true eigenvalues (computed via SVD).
Figure 5: GrEASE successfully approximates the SE of real-world datasets. A comparison between GrEASE and SpectralNet SE and Fiedler Vector (FV) approximation on real-world datasets. The values are the mean and standard deviation of the $sin^{2}$ distance between the predicted and true eigenvector of the test set, over 10 runs. Lower is better. GrEASE ability to separate the eigenvectors is evident.
Fig. 5 presents our results on the real-world datasets. GrEASE’s output is used directly, while SpectralNet’s predicted eigenvectors are resorted to minimize the mean $sin^{2}$ distance. The results clearly show that GrEASE consistently produces significantly more accurate SE approximations compared to SpectralNet, due to the improved separation of the eigenvectors.
Additionaly, note the GrEASE approximates the eigenvalues as well. When concerning a series of Laplacian eigenvalues, the most important property is the relative increase of the eigenvalues (Coifman and Lafon, 2006a). GrEASE demonstrates a strong ability to approximate this property. To see this, we repeated GrEASE’s eigenvalue approximation (10 times) and calculated the Pearson correlation between the predicted and accurate eigenvalues vector. We compared the first 10 eigenvalues. The resulting mean correlation and standard deviation are: Parkinsons Telemonitoring: $\textbf{0.917}_{\pm 0.0381}$; Appliances Energy Prediction: $\textbf{0.839}_{\pm 0.0342}$;

### v2

@@ 4.1 Motivation
It is well known that $V\in\mathbb{R}^{n\times k}$, whose columns are the first $k$ eigenvectors of $L$ (i.e., those corresponding to the $k$ smallest eigenvalues), minimizes $R_{L}(A)$ under orthogonality constraint (i.e., $A^{T}A=I$) (Li, 2015).
@@
An immediate result of Lemma 1 is that SpectralNet’s method, using a deep neural network for RQ minimization (while enforcing orthogonality), does not lead to the SE. However, it only leads to the space spanned by the constant vector and the leading k1k-1 eigenvectors of $L$, with different rotations and reflections for each run. Therefore, each time the RQ is minimized, it results in a different linear combination of the smallest eigenvectors. Although this is sufficient for clustering purposes, as we search for reproducibility, consistency, and separation of the eigenvectors, the RQ cannot solely provide the SE, necessitating the development of new techniques in Sep-SpectralNet.
4.2 Sep-SpectralNet
@@
Here we present the two key components of Sep-SpectralNet, a scalable and generalizable SE method. We consider the following setup: Given a training set $\mathcal{X}\subseteq\mathbb{R}^{d}$ and a target dimension $k$, we construct an affinity matrix $W$, and compute an approximation of the leading eigenvectors of its corresponding Laplacian. In practice, we first utilize SpectralNet (Shaham et al., 2018) to approximate the space spanned by the first $k+1$ eigenfunctions of the corresponding Laplace-Beltrami operator, and then find each of the $k$ leading eigenfunctions within this space (i.e. the SE). Namely, Sep-SpectralNet computes a map $F_{\theta}:\mathbb{R}^{d}\to\mathbb{R}^{k}$, which approximates the map $\bar{f}=(f_{2},\dots,f_{k+1})$, where $f_{i}$ is the $i$th eigenfunction of the Laplace-Beltrami operator on the underlying manifold $\mathcal{D}$.
@@
As empirically showed in (Shaham et al., 2018), and motivated from Lemma 1, SpectralNet loss is minimized when $F_{\theta}=T\circ(f_{1},\dots,f_{k+1})$, where $T:\mathbb{R}^{k+1}\to\mathbb{R}^{k+1}$ is an arbitrary isometry. That is, $F_{\theta}$ approximates the space spanned by the first $k+1$ eigenfunctions. However, the SE (i.e. each of the leading eigenfunctions) is poorly approximated. Each time the RQ is minimized, the eigenfunctions are approximated up to a different isometry $T$. Fig. 2(a) demonstrates this phenomenon on the toy moon dataset - a noisy half circle linearly embedded into 10-dimension input space (see Sec. 5.1). Employing SpectralNet approach indeed enables us to consistently achieve a perfect approximation of the space (i.e., the errors at the left histograms are accumulated around 0). However, when comparing vector to vector, it becomes apparent that the SE was seldom attained.
@@
To separate the eigenvectors, and thereby consistently get the SE, we present a simple use of Lemma 1. Notice that based on Lemma 1 we can compute a rotated version of the diagonal eigenvalues matrix. Namely,
$(VQ)^{T}L(VQ)=Q^{T}V^{T}LVQ=Q^{T}\Lambda Q=:\tilde{\Lambda}$,
where $\Lambda$ is the diagonal eigenvalues matrix. Due to the uniqueness of eigendecomposition, the eigenvectors and eigenvalues of the small matrix $\tilde{\Lambda}\in\mathbb{R}^{k+1\times k+1}$ are $Q^{T}$ and $\text{diag}(\Lambda)$, respectively. Hence, by diagonalizing $\tilde{\Lambda}$ we get the eigenvalues and are also able to separate the eigenvectors (i.e., approximate the SE).
@@
This step is simple and elegant, yet its impact is significant. The effect of this intentional rotation is represented in the Fig. 2(a). Sep-SpectralNet is not only able to consistently approximate the space, but also to approximate each eigenvector.
Figure 2: (a) Sep-SpectralNet separates the eigenvectors. Approximation of the 2-dimensional SE of the moon dataset (a 2D moon linearly embedded into 10D input space) using SpectralNet (in blue) and Sep-SpectralNet (in green) over 100 runs, on train (top row) and test (bottom row). Left column: distribution of the Grassmann distance between the output and true subspace. Second to Fourth columns: distribution of the $\text{sin}^{2}$ distance between each output and true eigenvector separately. Evidently, while SpectralNet’s errors are distributed over a large range of values, Sep-SpectralNet’s errors are small, capturing only the smallest error bin in the figure, indicating on eigenvector separation. (b) Sep-SpectralNet’s training is as scalable non-generalizable spectral methods. Running times of SE using Sep-SpectralNet vs. other methods on the moon dataset, relative to the number of samples, and with standard deviation confidence intervals. Evidently, Sep-SpectralNet is the fastest asymptotically.
@@
Our end-to-end training approach is summarized in Algorithms 1 and 2 in App. B. We first train $F_{\theta}$ to approximate the first eigenfunctions up to isometry (Algorithm 1) (Shaham et al., 2018), then compute $Q^{T}$ and $\Lambda$ to separate the eigenvectors and recover the SE and corresponding eigenvalues (Algorithm 2). App. C details additional considerations about the implementation.
@@
As the network iterates over small batches, and the post-processing operation is much cheaper, Sep-SpectralNet’s time complexity is approximately linear in the number of samples. This is also demonstrated in Fig. 2(b), where the continuous red line, representing linear regression, aligns with our empirical results. App. C provides a discussion about the complexity of Sep-SpectralNet. Sep-SpectralNet is also highly memory-efficient, requiring only minibatch-sized graphs or matrices in memory, rather than the full graph.
@@ 4.3 NUMAP
Here, we demonstrate Sep-SpectralNet’s extension to UMAP, one of many methods that may benefit from a generalizable SE. As discussed in Sec. 3.3, the SE initialization is crucial for the global preservation abilities of UMAP. Therefore, we seek a method to incorporate SE into a generalizable version of UMAP. A naïve approach would be to fine-tune Sep-SpectralNet using UMAP loss (Eq. 2). However, during this implementation, we encountered the phenomenon of catastrophic forgetting (see App. E).
The core of our idea is illustrated in Fig. 3. Initially, we use Sep-SpectralNet to learn a parametric representation of the $k$-dimensional SE of the input data. Subsequently, we train an NN to map from the SE to the final $\ell$-dimensional embedding space (usually for $\ell=2,3$), utilizing UMAP contrastive loss. The objective of the second NN is to identify representations that best capture the local structure of the input data graph. SE transforms complex non-linear structures into simpler linear structures, allowing the second NN to preserve both local and global structures effectively. To enhance this capability, we incorporate residual connections in the second NN, $G_{\theta}$, which link the first $\ell$ components of the input (the $\text{SE}_{1:\ell}$) directly to the output. Thus, the NUMAP embedding is given by $Y=\text{SE}(X)_{1:\ell}+G_{\theta}(X)$, that minimizes the UMAP’s loss
$\sum_{e\in E}w_{h}(e)\log\left(\frac{w_{h}(e)}{w_{l}(e)}\right)+(1-w_{h}(e))\log\left(\frac{1-w_{h}(e)}{1-w_{l}(e)}\right),$ (2)
where $w_{h}(e),w_{l}(e)$ are the corresponding weights of the edge $e$ in the high-dimensional input ($X$) graph and the low-dimensional output ($Y$) graph, respectively. It should be noted that this could not have been possible without Sep-SpectralNet’s ability to separate the eigenvectors, as the first $\ell$ components of SpectralNet’s output are merely arbitrary linear combinations of the $k$-dimensional SE. Also, it would not be practical without Sep-SpectralNet’s inherent generalizability and scalability. Fig. 1 demonstrates this capability with several simple structures.
Figure 3: Incorporating Sep-SpectralNet for generalizable UMAP. UMAP vs. NUMAP vs. P. UMAP overview. NUMAP integrates SE, as in UMAP, while enabling generalization. First, it learns a parametric representation of the $k$-dimensional SE. Then, it learns a map from the SE to the final $\ell$-dimensional space, incorporating residual connections from the first $\ell$ input components to the output layer.
@@ 4.4 Additional Applications
In this section we seek to highlight Sep-SpectralNet’s potential impact on important tasks and applications (besides generalizable UMAP), as it integrates generalizability, scalability and eigenvectors separation. As discussed in Sec. 1, SE is applied across various domains, many of which can benefit generalizability capabilities by replacing the current SE implementation with Sep-SpectralNet. We therefore elaborate herein the significance of SE in selected applications, and discuss how Sep-SpectralNet, as a generalizable approximation of it, can enhance their effectiveness and applicability.
@@
A special case of SE is the Fiedler vector and value (Fiedler, 1973; 1975). The Fiedler value, also known as algebraic connectivity, refers to the second eigenvalue of the Laplacian matrix, while the Fiedler vector refers to the associated eigenvector. This value quantifies the connectivity of a graph, increasing as the graph becomes more connected. Specifically, if a graph is not connected, its Fiedler value is 0. The Fiedler vector and value are a main topic of many works (Andersen et al., 2006; Barnard et al., 1993; Kundu et al., 2004; Shepherd et al., 2007; Cai et al., 2018; Zhu & Schlick, 2021; Tam & Dunson, 2020). As Sep-SpectralNet is able to distinguish between the eigenvectors and approximate the eigenvalues, it has the capability to approximate both the Fiedler vector and value, while also generalizing the vector to unseen samples (see Sec. 5.1).
@@
A popular method which incorporates SE, alongside the eigenvalues of the Laplacian matrix, is Diffusion Maps (Coifman & Lafon, 2006b). Diffusion Maps embeds a graph (or a manifold) into a space where the pairwise Euclidean distances are equivalent to the pairwise Diffusion distances on the graph. This approach is widely used (e.g., (Ghamari et al., 2024; Hardin et al., 2024; Park et al., 2021)).
@@
$X\to\begin{pmatrix}(1-\lambda_{2})^{t}v_{2}&\cdots&(1-\lambda_{k+1})^{t}v_{k+1}\end{pmatrix}=Y,$
where $X\in\mathbb{R}^{n\times d}$ is a matrix containing each input point as a row, and $Y\in\mathbb{R}^{n\times k}$ is a matrix containing each of the representations as a row. As Sep-SpectralNet is able to approximate both the eigenvectors and eigenvalues of the Laplacian matrix, it is able to make Diffusion Maps generalizable and efficient (Sec. 5.1).
@@ 4.5 Evaluating UMAP embedding - Grassmann Score
To address this gap, we propose a new evaluation method, specifically appropriate for assessing global structure preservation in graph-based dimensionality reduction methods (e.g., UMAP, t-SNE). The leading eigenvectors of the Laplacian matrix are known to encode crucial global information about the graph (Belkin & Niyogi, 2003). Thus, we measure the distance between the global structures of the original and embedding manifolds using the Grassmann distance between the first eigenvectors of their respective Laplacian matrices. We refer to this method as the Grassmann Score (GS).
It is important to note that GS includes a hyper-parameter - the number of eigenvectors considered. Increasing the number of eigenvectors incorporates more local structure into the evaluation. A natural choice for this hyperparameter is 2, which corresponds to comparing the Fiedler vectors (i.e., the second eigenvectors of the Laplacian). The Fiedler vector is well known for encapsulating the global information of a graph (Fiedler, 1973; 1975). Accordingly, for simplicity and unless specified otherwise, the GS is computed using the first two eigenvectors. Fig. 4 demonstrates GS (alongside Silhouette and kNN scores for comparison) on a few embeddings of two tangent spheres, independently to the embedding methods. Notably, the embedding on the right appears to best preserve the global structure, as indicated by the smallest GS value. In contrast, the kNN scores are comparable across all embeddings (e.g., kNN ignores separation of an existing class), and the Silhouette score even favors other embeddings. In App. D we mathematically formalize GS and provide additional examples of embeddings and their corresponding GS. These examples further support the intuition that GS effectively captures global structure preservation better than previous measures.
Figure 4: Grassmann Score (GS) reflects global structure preservation. A demonstration of the alignment between intuitive expectations and GS on a toy dataset of two 3D tangent spheres. Four possible 2D embeddings of the dataset are shown, alongside their corresponding GS, kNN accuracy, and Silhouette score. The rightmost visualization best preserves the global structure, forming two tangent circles in 2D. Notably, kNN fails to distinguish between the embeddings, while Silhouette favors the most clustered one, despite its poor reflection of the original structure. In contrast, GS successfully captures the preservation of global structure, assigning the highest score to the most faithful embedding.
@@ 5.1 Eigenvectors Separation - Generalizable SE
In this section, we demonstrate Sep-SpectralNet’s ability to approximate and generalize the SE using four real-world datasets: CIFAR10 (via their CLIP embedding); Appliances Energy Prediction dataset (Candanedo, 2017); Kuzushiji-MNIST (KMNIST) dataset (Clanuwat et al., 2018); Parkinsons Telemonitoring dataset (Tsanas & Little, 2009). Particularly, we compare our results with SpectralNet, which has been empirically shown to approximate the SE space. However, as our results demonstrate, SpectralNet is insufficient for accurately approximating SE. For additional technical details regarding the datasets, architectures and training procedures, we refer the reader to Appendix G.
@@
To assess the approximation of each eigenvector (i.e., the SE), we compute the $\text{sin}^{2}$ of the angle between each predicted and ground truth vector. This can be viewed as the 1-dimensional case of the Grassmann distance, a well-known metric for comparing equidimensional linear subspaces (see formalization in App. D). Concerning the eigenvalues approximation evaluation, we measure the Pearson Correlation between the predicted and true eigenvalues (computed via SVD).
Figure 5: Sep-SpectralNet successfully separates the eigenvectors. A comparison between Sep-SpectralNet and SpectralNet of SE and Fiedler Vector (v2) approximation on real-world datasets. The values are the mean and standard deviation of the $\text{sin}^{2}$ distance between the predicted and true eigenvector of the test set, over 10 runs. Lower is better. Notably, Sep-SpectralNet’s errors are significantly smaller than SpectralNet’s, supporting Sep-SpectralNet’s ability to separate the eigenvectors.
Fig. 5 presents our results on the real-world datasets. Sep-SpectralNet’s output is used directly, while SpectralNet’s predicted eigenvectors are resorted to minimize the mean $\text{sin}^{2}$ distance. The results clearly show that Sep-SpectralNet consistently produces significantly more accurate SE approximations compared to SpectralNet, due to the improved separation of the eigenvectors.
Additionaly, note the Sep-SpectralNet approximates the eigenvalues as well. When concerning a series of Laplacian eigenvalues, the most important property is the relative increase of the eigenvalues (Coifman & Lafon, 2006b). Sep-SpectralNet demonstrates a strong ability to approximate this property. To see this, we repeated Sep-SpectralNet’s eigenvalue approximation (10 times) and calculated the Pearson correlation between the predicted and accurate eigenvalues vector. We compared the first 10 eigenvalues. The resulting mean correlation and standard deviation are: Parkinsons Telemonitoring: $\textbf{0.917}_{\pm 0.0381}$; Appliances Energy Prediction: $\textbf{0.839}_{\pm 0.0342}$;